# Structure-Enhanced Protein Instruction Tuning: Towards General-Purpose Protein Understanding

### v2

@@ 5.1 Experimental Setting
In each stage of SEPIT, we utilize our proposed protein instruction dataset, employing different subsets at various stages. At Stage 0, our focus is primarily on basic protein descriptions derived from Swiss-Prot and the RCSB PDB. For Swiss-Prot, akin to previous work (Xu et al., 2023), we formulate the protein captions using functions, subcellular locations, and similarities. Regarding the RCSB PDB, we directly utilize abstracts from related PubMed papers collected by  (Guo et al., 2023) as captions. In Stage 1, we employee the same data as in Stage 0, but the output format is altered to the style of caption instructions. During Stage 2, we utilize the complete protein instruction dataset we proposed, which includes open-ended generation and closed-set answers tasks.
@@
We evaluate the capability of SEPIT for general-purpose protein understanding on the test set of the protein instruction dataset we proposed with the state-of-the-art models. There are four main categories of methods. Among the Zero-Shot methods, we include current mainstream LLMs providing API services (e.g., Claude-3-haiku (Anthropic, 2024), GPT-3.5-turbo (OpenAI, 2024a), and GPT-4-turbo (OpenAI, 2024b)), open-source LLMs fine-tuned on biomedical corpus (e.g., Galactica (Taylor et al., 2022), BioMedGPT (Luo et al., 2023b)) and open-source LLMs fine-tuned specifically on molecular or protein knowledge(e.g., Mol-Instructions (Fang et al., 2024), BioT5+ (Pei et al., 2024a) and InstructProtein (Wang et al., 2023b)). In the category of instruction tuning methods, we evaluate mainstream open-source LLMs (e.g., TinyLlama-Chat (Zhang et al., 2024), OpenLlama-v2 (OpenLLMAI, 2023), Llama2-Chat (Touvron et al., 2023)), where the protein sequences are input in natural language form. For sequence-only protein instruction tuning methods (PITs), ESM2-650M (Lin et al., 2023) was utilized as the protein encoder, with only protein sequences input. In addition to our proposed SEPIT framework, SEPIT-TinyLlama-MoEs, we have also designed two variants that differ in the LLMs’ architecture. For evaluation metrics, we employ BLEU score (Sutskever et al., 2014), ROUGE score (Lin, 2004), METEOR score (Banerjee & Lavie, 2005), BERT score (Zhang et al., 2019) calculated by PubMedBERT (Gu et al., 2021), and Accuracy to assess performance across two types of tasks: open-ended generation and closed-set answer, respectively. It is worth noting that related works (Lv et al., 2024; Guo et al., 2023; Liu et al., 2024b) exist that can caption protein sequences but lack instruction-following abilities. Direct comparison with these methods is not possible here. However, we have compared representative methods of them using an alternative approach detailed in the Appendix B.3.
5.2 Performance Comparisons
@@
Table 1: Performance comparisons on open-ended generation and closed-set answer tasks.
Model Activated Parameters Open-ended Closed-set BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L METEOR BERT-P BERT-R BERT-F1 Accuracy Zero-Shot GPT-3.5-turbo N/A 3.26 0.02 12.41 3.14 11.06 10.44 85.18 85.40 85.24 56.56% Claude-3-haiku N/A 3.00 0.07 12.10 2.65 10.62 9.28 86.04 85.47 85.70 59.14% GPT-4-turbo N/A 4.21 0.08 12.78 2.93 11.57 10.41 86.91 85.56 85.71 58.58% Galactica 1.3B 0.43 0.01 3.49 0.41 2.67 2.44 85.79 82.61 84.08 39.15% BioMedGPT 7B 0.83 0.01 4.90 0.49 3.26 4.59 85.51 84.95 85.14 38.61% Mol-Instructions 7B 0.53 0.01 5.96 0.39 4.64 5.51 83.81 84.41 84.06 BioT5+ 252M 3.88 1.92 12.12 4.88 10.37 14.26 85.14 85.93 85.48 InstructProtein 1.3B 5.50 2.97 14.80 5.68 13.76 13.17 85.34 85.92 85.57 48.37% Instruction Tuning TinyLlama 1.1B 51.16 43.44 65.41 51.26 62.31 60.80 93.97 94.37 94.16 74.09% OpenLlama-v2 3B 36.19 30.65 48.33 36.52 45.53 49.01 92.92 91.87 92.35 71.77% Llama2 7B 57.02 49.47 70.80 57.24 67.78 65.96 94.95 95.17 95.05 71.68% Sequence-Only Protein Instruction Tuning PIT-TinyLlama 1.8B 57.82 50.02 71.34 58.16 68.35 66.19 95.18 95.28 95.26 76.02% PIT-TinyLlama-MoEs 1.8B 57.92 50.01 72.13 58.21 69.19 66.29 95.31 95.30 95.29 78.56% Structure-Enhanced Protein Instruction Tuning SEPIT-TinyLlama 1.8B 58.43 51.04 72.34 58.77 69.13 67.91 95.32 95.59 95.44 79.05% SEPIT-Llama2 8B 60.81 52.37 74.80 60.84 71.62 68.43 95.81 95.73 95.76 79.97% SEPIT-TinyLlama-MoEs 1.8B 60.28 52.16 74.22 60.29 71.13 68.27 95.62 95.69 95.64 79.73%
5.3 Ablation Study
@@
For SEPIT’s model architecture, we propose the following variants: without the
Table 2: Ablation study on SEPIT’s architecture.
Table 3: Performance of SEPIT’s encoder.
structure-aware module (w/o Structure), without the mixture of experts module (w/o MoEs), without both the aforementioned modules (w/o Structure & MoEs), without Stage 0 pre-training (w/o Stage 0), and completely excluding the SEPIT framework (w/o SEPIT). Table 3 shows the results of the ablation study, proving the significant effectiveness of each component within SEPIT’s model architecture. The absence of either the structure-aware module or the MoEs Module leads to performance degradation in both open-ended generation tasks and closed-set answer tasks, with further deterioration when both are excluded. Meanwhile, the performance under w/o SEPIT intuitively demonstrates the overall effectiveness of the framework. It is noteworthy that the results for w/o Stage 0 are not available, as under the implementation using automatic mixed precision (AMP) based on FP16, the randomly initialized structure-aware module would bring excessively large gradients causing overflow, even though we employed a warm-up learning rate scheduler. Due to device restrictions, we were unable to use BF16 type; however, this issue was resolved as Stage 0 progressed. In order to supplement the analysis of the effectiveness of Stage 0, we validate the protein encoder pre-trained by Stage 0 on commonly used EC, GO annotation tasks. (Gligorijević et al., 2021). The results, as shown in Table 3, demonstrates the performance comparisons to state-of-the-art methods on $\text{F}_{\text{max}}$.
@@
Regarding the data, apart from using Swiss-Prot and RCSB PDB for constructing the protein instruction dataset, there exists a substantial amount of protein-text paired data in TrEMBL (Bairoch & Apweiler, 1997). Considering that the TrEMBL data is annotated by automated methods and has not been manually screened, we select proteins with more comprehensive descriptions (annotation score $\geq 4$) to construct a supplementary dataset using the same method, with a sample size (5.25M) close to the entire protein instruction dataset (5.47M). Disappointingly, as shown in Table 3 (w/ TrEMBL), even after doubling the GPUs cost, what we obtain is a decrease in performance. Our analysis suggests that directly mixing low-quality data into high-quality data introduces noise, and protein understanding tasks require higher quality over quantity for data. More results can be found in Appendix B.4.
Figure 3: The workload of experts in SEPIT (left) and tokens’ pathways among experts (right).
5.4 Case Study
Consistency of SEPIT with Different Protein Input Formats.
@@ 5.4 Case Study
Table 4: Performence of SEPIT with different protein input formats.
@@
Model Train w/ Struct. Infer w/ Struct. Open-ended Generation Closed-set Answer BLEU-2 ROUGE-L METEOR BERT-F1 Accuracy PIT-TinyLlama 57.82 68.35 66.19 95.26 76.02% PIT-TinyLlama-MoEs 57.92 69.19 66.29 95.29 78.56% SEPIT-TinyLlama 58.43 69.13 67.91 95.44 79.05% 57.95 68.75 67.54 95.38 77.80% SEPIT-Llama 60.81 71.62 68.43 95.76 79.97% 60.64 71.48 68.34 95.74 79.91% SEPIT-TinyLlama-MoEs 60.28 71.13 68.27 95.64 79.73% 59.98 70.87 68.00 95.59 79.53%
@@
Table 5: Case studies on general-purpose protein understanding ability of SEPIT.
@@
[Uncaptioned image] P9WF36 (Swiss-Prot) Question: Explain the function of <protein>. Ground Truth: Acts as a transcriptional regulator. Probably redox-responsive. The apo- but not holo-form probably binds DNA (By similarity). SEPIT-TinyLlama-MoEs: Acts as a transcriptional regulator. Probably redox-responsive. The apo- but not holo-form probably binds DNA. PIT-TinyLlama-MoEs: Acts as a transcriptional regulator, probably involved in antibiotic resistance gene expression. Binds DNA (By similarity). Llama-Chat: Involved in resistance toward heavy metals (PubMed:12839546). Binds cadmium or lead ions but not zinc ions (By similarity). May be involved in intracellular transport of Cd(+) or Pb(+), as well as their detoxification by sequestration into vacuoles (Probable) (PubMed:17081065). GPT-4-turbo: The string you’ve provided appears to be an amino acid sequence of a protein. However, without additional context or data, I cannot provide a function caption for this specific sequence directly. [Uncaptioned image] A0KR49 (Swiss-Prot) Question: Which GO terms outline the biological processes of <protein>? Ground Truth: fatty acid beta-oxidation SEPIT-TinyLlama-MoEs: fatty acid beta-oxidation PIT-TinyLlama-MoEs: fatty acid beta-oxidation; phenylacetate catabolic process Llama-Chat: fatty acid beta-oxidation; phenylacetate catabolic process GPT-4-turbo: To understand the biological processes of the provided protein sequence, one must first consider its origin, functions, and structure. However, without direct access to databases or running bioinformatic tools right now…
Workload of Experts in SEPIT.
In SEPIT, we utilize mixture of experts, and in Figure 3, we present the workload distribution of different experts during inference on test set of the open-ended generation task. In the left graph, we can observe that the experts are evenly activated, indicating that the auxiliary loss has played its expected role, which lays the foundation for efficient parallel inference of experts. In the right graph, we visualize the pathways distribution of text and protein tokens across experts in different layers, and we have observed an intriguing phenomenon. Unlike the findings in previous vision-language multimodal research (Lin et al., 2024), in SEPIT, text and protein tokens are processed by different experts instead of following almost identical pathways as do text and image tokens. We believe that this stems from the fundamental difference between proteins and images. That is, protein tokens, which represent amino acids, cannot reflect the properties of the entire protein, while image tokens represent specific regions of an image containing independent information that can correspond to a part of the image’s caption. This validates our choice to use complete protein representation sequences as inputs for LLMs, rather than compressing tokens as is often done in vision-language tasks. More visualization is shown in Appendix B.5.
General-Purpose Protein Understanding Ability of SEPIT.
@@
At last, we would like to showcase the general-purpose protein understanding capabilities of SEPIT. As shown in Table 5, we present two cases from the protein instruction dataset’s test. For case 1, SEPIT accurately responds regarding the protein’s function, whereas PIT incorporates incorrect details, and both Llama-Chat and GPT-4 offer entirely inaccurate responses. For case 2, SEPIT also gives the correct response, while the answers from PIT and Llama-Chat, although covering the correct answer, come with additional incorrect information, likely due to hallucinations caused by the lack of structural information. Due to space limitations, more cases are included in the Appendix B.5.
6 Conclusion
@@
In this work, we introduce SEPIT, a novel approach for general-purpose protein understanding. SEPIT aims to enable LLMs to interpret both the sequence and structural information of proteins, thus following instructions to generate specific understanding on protein properties and functions. To achieve this, we integrate structure-aware enhancements into pre-trained pLM and connect them to LLMs via a linear projector. The models are then trained in a two-stage instruction tuning pipeline on protein instruction dataset we constructed which is the largest protein instruction dataset to date. Experimental results show that SEPIT significantly outperforms the state-of-the-art models.
Appendix
Appendix A Construction of Protein Instruction Dataset

### v3

@@ 5.1. Experimental Setting
In each stage of SEPIT, we utilize our proposed protein instruction dataset, employing different subsets at various stages. At Stage 0, our focus is primarily on basic protein descriptions derived from Swiss-Prot and the RCSB PDB. For Swiss-Prot, akin to previous work (Xu et al., 2023), we formulate the protein captions using Function, Similarity and Subcellular location. Regarding the RCSB PDB, we directly utilize abstracts from related PubMed papers collected by  (Guo et al., 2023) as captions. In Stage 1, we employ the same data as in Stage 0, but the output format is altered to the style of caption instructions. During Stage 2, we utilize the complete protein instruction dataset we proposed, which includes open-ended generation and closed-set answers tasks.
@@
We evaluate the capability of SEPIT for general-purpose protein understanding on the test set of the protein instruction dataset we proposed with the state-of-the-art models. There are four main categories of methods. Among the Zero-Shot methods, we include current mainstream LLMs providing API services (e.g., Claude-3-haiku (Anthropic, 2024), GPT-3.5-turbo/4-turbo/4o-mini/4o (OpenAI et al., 2024) and DeepSeek-V3 (DeepSeek-AI et al., 2025)), open-source LLMs fine-tuned on biomedical corpus (e.g., Galactica (Taylor et al., 2022), BioMedGPT (Luo et al., 2023b)) and open-source LLMs fine-tuned specifically on molecular or protein knowledge(e.g., Mol-Instructions (Fang et al., 2024), BioT5+ (Pei et al., 2024a), InstructProtein (Wang et al., 2024c) and ProtT3 (Liu et al., 2024b)). In the category of instruction tuning methods, we evaluate mainstream open-source LLMs (e.g., TinyLlama-Chat (Zhang et al., 2024), OpenLlama-v2 (Geng and Liu, 2023), Llama2-Chat (Touvron et al., 2023)), where the protein sequences are input in natural language form. For sequence-only protein instruction tuning methods (PITs), ESM2-650M (Lin et al., 2023) was utilized as the protein encoder, with only protein sequences input. In addition to our proposed SEPIT framework, SEPIT-TinyLlama-MoEs, we have also designed two variants that differ in the LLMs’ architecture. For evaluation metrics, we employ BLEU score (Sutskever et al., 2014), ROUGE score (Lin, 2004), METEOR score (Banerjee and Lavie, 2005), BERT score (Zhang* et al., 2020) calculated by PubMedBERT (Gu et al., 2021), and Accuracy to assess performance across two types of tasks: open-ended generation and closed-set answer, respectively.
Table 3. Ablation study on SEPIT’s architecture.
5.2. Performance Comparisons
@@
Table 4. Performance of SEPIT’s encoder.
Considering that there are additional related works (Lv et al., 2024; Guo et al., 2023; Liu et al., 2024b) capable of captioning input protein sequences but lacking instruction-following capabilities, we attempted to compare performance on overlapping properties and functions with these works. Table 2 shows a performance comparison with ProtT3 (Liu et al., 2024b), which is the most representative among these works, demonstrates superior performance. For open-ended generation and closed-set answer tasks, we specifically selected Function, Similarity, Subcellular location, and overlapping Q&A about proteins from RCSB PDB for comparison. The results demonstrate that our model exhibits significantly better performance.
5.3. Ablation Studies
@@
For SEPIT’s model architecture, we propose the following variants: without the structure-aware module (w/o Structure), without the mixture of experts module (w/o MoEs), without both the aforementioned modules (w/o Structure & MoEs), without Stage 0 pre-training (w/o Stage 0), and completely excluding the SEPIT framework (w/o SEPIT). Table 3 shows the results of the ablation study, proving the significant effectiveness of each component within SEPIT’s model architecture. The absence of either the structure-aware module or the MoEs Module leads to performance degradation in both open-ended generation tasks and closed-set answer tasks, with further deterioration when both are excluded. Meanwhile, the performance under w/o SEPIT intuitively demonstrates the overall effectiveness of the framework. It is noteworthy that the results for w/o Stage 0 are not available, as under the implementation using automatic mixed precision (AMP) based on FP16, the randomly initialized structure-aware module would bring excessively large gradients causing overflow, even though we employed a warm-up learning rate scheduler. Due to device restrictions, we were unable to use BF16 type; however, this issue was resolved as Stage 0 progressed. In order to supplement the analysis of the effectiveness of Stage 0, we validate the protein encoder trained by Stage 0 on commonly used EC, GO annotation tasks (Gligorijević et al., 2021). The results, as shown in Table 4, demonstrate the performance comparisons to state-of-the-art methods on $\text{F}_{\text{max}}$ metric.
@@
Regarding the data, apart from using Swiss-Prot and RCSB PDB for constructing the protein instruction dataset, there exists a substantial amount of protein-text paired data in TrEMBL (Bairoch and Apweiler, 1997). Considering that the TrEMBL data is annotated by automated methods and has not been manually screened, we select proteins with more comprehensive descriptions (annotation score $\geq 4$) to construct a supplementary dataset using the same method, with a sample size (5.25M) close to the entire protein instruction dataset (5.47M). Disappointingly, as shown in Table 3 (w/ TrEMBL), even after doubling the GPUs cost, what we obtain is a decrease in performance. Our analysis suggests that directly mixing low-quality data into high-quality data introduces noise, and protein understanding tasks require higher quality over quantity for data. More results can be found in Appendix C.3.
Table 5. Performance comparisons with different protein input formats (all models are based on TinyLlama).
Model PIT PIT SEPIT SEPIT SEPIT SEPIT -MoEs -MoEs -MoEs Input Formation Train w/ Struct. Infer w/ Struct. Open-Ended BLEU-2 57.82 57.92 58.43 57.95 60.28 59.98 ROUGE-L 68.35 69.19 69.13 68.75 71.13 70.87 METEOR 66.19 66.29 67.91 67.54 68.27 68.00 BERT-F1 95.26 95.29 95.44 95.38 95.64 95.59 Closed-Set Accuracy 76.02% 78.56% 79.05% 77.80% 79.73% 79.53%
5.4. Generalization Analysis
Generalization Across Different Protein Input Formats.
@@ 5.4. Generalization Analysis
Table 6. Performance comparisons on OOD proteins.
@@
Model BLEU-2 ROUGE-L METEOR BERT-F1 TinyLlama 22.23 35.39 35.27 90.26 Llama2 28.51 41.63 41.58 91.27 PIT-TinyLlama 37.59 50.60 48.61 92.73 PIT-TinyLlama-MoEs 38.37 50.09 51.10 92.70 SEPIT-TinyLlama 40.25 52.03 52.18 93.02 + w/o input structure 39.75 51.84 52.24 93.00 SEPIT-TinyLlama-MoEs 41.60 54.26 51.91 93.30 + w/o input structure 41.26 53.85 51.52 93.23
Generalization to Out-of-Distribution Proteins.
To further evaluate the generalization ability of SEPIT on highly novel proteins, we constructed an out-of-distribution (OOD) test dataset by filtering sequences with a maximum amino acid similarity of $<40\%$ using the CD-HIT clustering algorithm (Fu et al., 2012). As shown in Table 6, all methods exhibited performance drops on this challenging setup, with instruction tuning methods being particularly affected. However, SEPIT demonstrated stronger generalization compared to alternative approaches, owing to its integration of both pLMs’ prior knowledge and the structural information. Remarkably, SEPIT maintained its performance even without structural input during inference, highlighting its robust generalization capacity.
@@
Table 7. Case studies on general-purpose protein understanding ability of SEPIT.
@@
[Uncaptioned image] P9WF36 (Swiss-Prot) Question: Explain the function of ¡protein¿. Ground Truth: Acts as a transcriptional regulator. Probably redox-responsive. The apo- but not holo-form probably binds DNA (By similarity). SEPIT-TinyLlama-MoEs: Acts as a transcriptional regulator. Probably redox-responsive. The apo- but not holo-form probably binds DNA. PIT-TinyLlama-MoEs: Acts as a transcriptional regulator, probably involved in antibiotic resistance gene expression. Binds DNA (By similarity). Llama-Chat: Involved in resistance toward heavy metals (PubMed:12839546). Binds cadmium or lead ions but not zinc ions (By similarity). May be involved in intracellular transport of Cd(+) or Pb(+), as well as their detoxification by sequestration into vacuoles (Probable) (PubMed:17081065). GPT-4-turbo: The string you’ve provided appears to be an amino acid sequence of a protein. However, without additional context or data, I cannot provide a function caption for this specific sequence directly. [Uncaptioned image] A0KR49 (Swiss-Prot) Question: Which GO terms outline the biological processes of ¡protein¿? Ground Truth: fatty acid beta-oxidation SEPIT-TinyLlama-MoEs: fatty acid beta-oxidation PIT-TinyLlama-MoEs: fatty acid beta-oxidation; phenylacetate catabolic process Llama-Chat: fatty acid beta-oxidation; phenylacetate catabolic process GPT-4-turbo: To understand the biological processes of the provided protein sequence, one must first consider its origin, functions, and structure. However, without direct access to databases or running bioinformatic tools right now… [Uncaptioned image] Q7CS13 (Swiss-Prot) Question: What is the protein family of ¡protein¿? Ground Truth: Belongs to the metallo-dependent hydrolases superfamily. Atu3266/EF_0837 deacetylase family. SEPIT-TinyLlama-MoEs: Belongs to the metallo-dependent hydrolases superfamily. PIT-TinyLlama-MoEs: Belongs to the metallo-dependent hydrolases superfamily. Type II family. Llama-Chat: Belongs to the metallo-dependent hydrolases superfamily. HutI family. GPT-4-turbo: To understand the biological processes of the provided protein sequence, one must first consider its origin, functions, and structure. However, without direct access to databases or running bioinformatic tools right now…
Figure 3. The workload of experts in SEPIT (left) and tokens’ pathways among experts (right).
5.5. Case Studies
General-Purpose Protein Understanding Ability of SEPIT
@@ 5.5. Case Studies
At last, we would like to showcase the general-purpose protein understanding capability of SEPIT. As shown in Table 7, we present three cases from the test set of our protein instruction dataset. For case 1, SEPIT accurately responds regarding the protein’s function, whereas PIT incorporates incorrect details, and both Llama-Chat and GPT-4 offer entirely inaccurate responses. For case 2, SEPIT also gives the correct response, while the answers from PIT and Llama-Chat, although covering the correct answer, come with additional incorrect information, likely due to hallucinations caused by the lack of structural information. For case 3, SEPIT provides the partially correct response, whereas PIT and Llama-Chat exhibit a similar error pattern to case 2. Due to space limitations, more cases and analysis are included in the Appendix C.4, C.5.
Workload of Experts in SEPIT
In SEPIT, we utilize mixture of experts, and in Figure 3, we present the workload distribution of different experts during inference on test set of the open-ended generation task. In the left graph, we can observe that the experts are evenly activated, indicating that the auxiliary loss has played its expected role, which lays the foundation for efficient parallel inference of experts. In the right graph, we visualize the pathways distribution of text and protein tokens across experts in different layers, and we observe an intriguing phenomenon. Unlike the findings in previous vision-language multimodal research (Lin et al., 2024), in SEPIT, text and protein tokens are processed by different experts instead of following almost identical pathways as do text and image tokens. We believe that this stems from the fundamental difference between proteins and images. That is, protein tokens, which represent amino acids, cannot reflect the properties of the entire protein, while image tokens represent specific regions of an image containing independent information that can correspond to a part of the image’s caption. This validates our choice to use complete protein representation sequences as inputs for LLMs, rather than compressing tokens as is often done in vision-language tasks. More visualization is shown in Appendix C.4.
6. Conclusion
@@
In this work, we introduce SEPIT, a novel approach for general-purpose protein understanding. SEPIT is designed to enable LLMs to interpret both the sequence and structural information of proteins, allowing them to follow instructions and generate precise insights into protein properties and functions. To achieve this, we integrate structure-aware enhancements into pre-trained pLMs and connect them to LLMs via a linear projector. The models are then trained using a two-stage instruction tuning pipeline on the largest protein instruction dataset to date, which we constructed. Experimental results demonstrate that SEPIT significantly outperforms the state-of-the-art models, highlighting its effectiveness in advancing general-purpose protein understanding.
Acknowledgment
This work was supported in part by the National Key R&D Program of China (Grant No.2023YFF0725001), in part by the National Natural Science Foundation of China (Grant No.92370204), in part by the Guangdong Basic and Applied Basic Research Foundation (Grant No.2023B1515120057), Education Bureau of Guangzhou Municipality.
Appendix A Supplement to Related Work
Multimodal Instruction Tuning.
With the emergence of Multimodal LLMs (MLLMs) such as GPT4 (OpenAI et al., 2024) and Genimi (Team et al., 2025), MLLMs have become a focal point of research. Initially, works like CLIP (Radford et al., 2021a), ALBEF (Li et al., 2021), VLMo (Radford et al., 2021b), SimVLM (Wang et al., 2021) emphasized exploring the cross-modal alignment (Chen et al., 2024a) between vision and language. Subsequently, based on modal alignment, Flamingo (Alayrac et al., 2022) and BLIP2 (Li et al., 2023) established bridges between visual encoders and LLMs using the Perceiver Resampler and the Q-Former, respectively. Following this, PaLM-E (Driess et al., 2023) introduced ”multimodal sentences” as input, injecting real-world continuous sensor data into the LLMs in the form of language tokens, thereby endowing the model with a general multi-task capability. Additionally, efforts such as InstructBLIP (Dai et al., 2024), LLaVA (Liu et al., 2024a), MiniGPT4 (Zhu et al., 2024), mPLUGOwl (Ye et al., 2024), Qwen-VL (Bai et al., 2023), CogVLM (Wang et al., 2024b) applied the crucial instruction tuning technique from LLMs to MLLMs, enhancing the MLLMs’ ability to follow multimodal instructions. At the same time, they introduced innovations from the perspectives such as the instruction construction, training paradigms, and model design, which in turn refreshed the performance of MLLMs in a variety of visual-language downstream tasks (Yin et al., 2024). In this paper, we attempt to apply this paradigm to the protein domain, investigating the potential of endowing LLMs with general-purpose protein understanding capabilities.
Learning with 3D Structural Information.
Although pLMs pre-trained on protein sequences have been proven to be effective in numerous tasks, the protein structure is inherently a determinant of protein function (Anfinsen and Haber, 1961; Hartley, 1951). More effectively utilizing 3D structural information can help to understand proteins more comprehensively. To capture the impact of geometric positioning and interaction relationships among residues, a class of methods that encode 3D geometric information into rotation-invariant scalars, which was then processed through graph neural networks (GNNs) (Wu et al., 2020, 2024; Xiao et al., 2023; Chen et al., 2024b) for message passing (Wu et al., 2022; Zhang et al., 2023a). For example, IEConv(Hermosilla et al., 2021) utilized a multi-graph to depict primary and secondary structures through covalent and hydrogen bonds and represented the tertiary structure with the spatial 3D coordinates of atoms. By blending intrinsic and extrinsic node distances and employing hierarchical pooling, it effectively perceived all three structural levels of proteins. GearNet went further by incorporating three types of directed edges (sequential edges, radius edges and k-NN edges) into the graph, capturing information at various structural levels. On this basis, CDConv (Fan et al., 2023) attempted to parameterize the kernel matrices using MLP, as opposed to employing distinct kernel matrices for varying edge types, which enabled a more flexible and efficient modeling of complex interactions between residues. Additionally, another class of methods sought to prevent the loss of 3D structural information by incorporating 3D rigid transformations into the network operations. This led to the development of geometric GNNs/Transformers characterized by SE(3) invariance and equivariance (Liao and Smidt, 2022; Liao et al., 2024; Fuchs et al., 2020; Zhou et al., 2023; Luo et al., 2023a; Wang et al., 2024a). Representative examples of this approach, such as GVP (Jing et al., 2021) and EvoFormer (Jumper et al., 2021), were utilized by ESM-IF (Hsu et al., 2022), AlphaFold2 (Jumper et al., 2021), respectively. Furthermore, to leverage information from both evolutionary-scale protein sequences and the relatively limited protein structures, some other methods (Wang et al., 2022; Zhang et al., 2023a) attempted to establish a connection between these two types of data.
Appendix B Construction of Protein Instruction Dataset